v_{f} = v_{i} t

ω_{f} = ω_{i} t

F⃗ = m a⃗

τ⃗ = I α⃗

F⃗ = \frac{dp⃗}{dt}

τ⃗ = \frac{dL⃗}{dt}

p⃗ = m v⃗

L⃗ = I ω⃗

L⃗ = I ω⃗ (just like p⃗ = m v⃗)

L⃗ = r⃗ × p⃗

\frac{dp⃗}{dt} = m\frac{dv⃗}{dt} = ma⃗ = F⃗

\frac{dL⃗}{dt} = I \frac{dω⃗}{dt} = Iα⃗ = τ⃗

|L| = |r| |p| sin(θ)